OpenStudy (madeleine2):

Write the standard form of the line that is parallel to 2x + 3y =4 and passes through the point (1, -4).

OpenStudy (3mar):

May I help?

OpenStudy (madeleine2):

yes @3mar

OpenStudy (3mar):

Thanks. This would help you. Tell me how do you plug in your givings!

OpenStudy (madeleine2):

i dont get it

OpenStudy (madeleine2):

@3mar

OpenStudy (3mar):

Yes, I am here

OpenStudy (3mar):

parallel to 2x + 3y =4 means that the slope of our line is the same as the slope of this line 2x+3y=4 3y=4-2x y=(4/3)-(2/3)x so the slope will be 2/3

OpenStudy (3mar):

You have the slope and the point now, so what are you waiting for?

OpenStudy (3mar):

@madeleine2